OpenStudy (marsxtc):

Find the points of intersection. y=x^3-3 and y=5-2x I tried doing x^3-3=5-2x But Im really not sure what to do.. Help!

OpenStudy (anonymous):

[x^3-3=5-2x x^3-2x-3=0 now you can solve for x or by graphing this